What is the exact area of a circle with a diameter of 31 inches?
Answer by stanbon(75887)

What is the exact area of a circle with a diameter of 31 inches?
---
Area = pi*r^2
----
Area = (pi)(15.5)^2 = 240.25(pi) sq in
================
Cheers,
Stan H.